# Generative Physical AI in Vision: A Survey

### v1

@@ 6 Physics Evaluation
Most of the image and video generation methods are evaluated based on FID [193], CLIPSIM [194], CLIP-FID [195], and FVD [196] to evaluate visual quality and diversity.
However, there are a number of gaps in these metrics. Frechet Inception distance (FID) [193], Clip Similarity (CLIPSIM) [194], CLIP-FID [195] and Inception score (IS) [193] are more biased toward the alignment between images’ contents and text prompt. For video generation, the detection of unrealistic motions violating physical laws is more important, which cannot be detected by these metrics effectively. Furthermore, the evaluation of FVD requires a reference video dataset, which is difficult to have for newly generated scenes. These challenges prevent these metrics from effectively evaluating physical common sense. Some works use Vision Language Models (VLMs) as evaluators [197, 198], yet these works only evaluate the spatial relationships. Most of them face challenges when generalizing to physical correctness.
In this section, we introduce the benchmarks and metrics designed for evaluating the physics awareness of image or video generation models.
@@ 6.1 Benchmarks
Category Physical laws
Mechanics gravity, buoyancy, solid pressure, atmospheric pressure,
elasticity, friction, surface tension
Optics reflection, refraction, scattering, dispersion,
interference and diffraction, straight-line propagation
Thermal Solidification, Melting, Liquefaction, Boiling,
deposition, Sublimation
Material color, hardness, solubility, combustibility and flame reaction,
acidity, redox potential, dehydrating properties
TABLE III: Taxonomy of physical commonsense in PhyGenBench[199].
@@
PhyBench. PhyBench[200] and PhyGenBench[199] define physical commonsense as four types, including mechanics, optics, thermal and material properties associated with physical laws as in Table III. PhyBench [200] constructs the benchmark by extending the designed implicit prompt describing a type of physical commonsense. For example, given “pressure" topic, the authors design a prompt “an empty plastic bottle at the bottom of the sea". This is because the bottle would stand a very large pressure when staying under the water. After that, “empty plastic bottle" will be replaced by other objects, and the scenario "at the bottom of the sea" is also replaced by other scenarios to describe other topics. This replacement process is conducted using GPT-4.
PhyGenBench. Due to the limited scope of PhyBench, PhyGenBench [199] constructs the benchmark with 4 categories from Table III for video generation through five steps:
  • Conceptualization: Identify key physical commonsense from four types of physical commonsense
  • Prompt Engineering: Craft initial prompts to describe the underlying physical phenomenon.
  • Prompt Augmentation: Adding additional details to prompt
  • Diversity Enhancement: Use GPT-4o to do object substitution on the augmented prompts
  • Quality control: Review the prompts and associated physical laws for accuracy and relevance.
VideoPhy. Besides the four physical categories defined in Table III, VideoPhy[201] categorizes the physical interaction types into three categories which are Solid-Solid, Solid-Fluid and Fluid-Fluid. An example of Solid-Solid is the prompt “A cat is playing with a bottle," where the interaction between “cat" and “bottle" is solid-solid, or “Water flows in an open pipeline" is the interaction between “water" and “pipeline" describing Solid-Fluid. Similarly, the prompt “Rain is dropping onto a river" is a Fluid-Fluid interaction. Based on these definitions, VideoPhy [201] constructs a total of 688 captions for the benchmark.
@@ 6.2 Evaluation metrics
There are two main approaches for the evaluation metrics of physical awareness. The first approach is based on human evaluation, where humans will give the score to the image or video in terms of physical commonsense with defined criteria. The second way is automatic evaluation.
For human evaluation, this method achieves very high expectations for the evaluation process [200, 199, 201, 202]. However, this results in exorbitant cost due to the involvement of humans, which is also slower.
For automatic evaluation, there are two approaches, i.e., obtaining scores or yes/no answers from LVMs/LLMs models [200, 199, 201] or obtaining specialized quantitative scores.
Specialized Quantitative Scores. In [192], the authors used Stream Function Error (SFE) [203, 204] for assessing streamline accuracy, Smoothness Error (SE) [205] for measuring stable propagation characteristics. Gradient Smoothness (GS) and Continuity Score (CS) [204] are used to evaluate the reasonableness and consistency of the phenomenon without using ground truth. Q-Criterion Error (QCE) [206] and Vorticity Error (VE) [207] are used to measure the preservation of vortex structure and vorticity accuracy accordingly.
LVMs-Based Scores. There are two approaches by which we can utilize the LVMs model to evaluate the physical commonsense. The first method is to ask the LVM model to give a score for scene and physical correctness assessments directly according to some criteria [200]. The model is also required to provide justification for its score. In [199], the author further improves the evaluation by evaluating three different scores, which are “Key Physical Phenomena Detection", “Physics Order Verification," and “Overall Naturalness Evaluation". The scores obtained from these three criteria will be averaged to achieve the final score.
  • Key Physical Phenomena Detection: This score is to decide if the physical phenomena occur in the video.
  • Physics Order Verification: This score will assess if the physical phenomena occurred in the correct order.
  • Overall Naturalness Evaluation: This score aims to assess the overall naturalness of the video.
Besides from asking the score directly from the LVMs, [201] asks the LVM to answer if the generated video/image has satisfied Semantic Adherence (SA) or Physical Commonsense (PC). The answer from LVM is required to be “Yes" or “No". The final score will be calculated as:
$s_{\theta}(\tau_{t}(x))=\frac{p_{\theta}(\text{Yes}|\tau_{t}(x))}{p_{\theta}(\text{Yes}|\tau_{t}(x))+p_{\theta}(\text{No}|\tau_{t}(x))}$ (14)
where $p_{\theta}(\text{Yes}|\tau_{t}(x))$ is the probability that the LVM model will answer “Yes" with $t\in\{SA,PC\}$.
@@ 7 Discussions
Semantic Understanding vs. Physical Understanding. Semantic understanding and physics understanding serve distinct but complementary roles in how models interpret and generate visual data. Semantic understanding involves interpreting and recognizing concepts and objects within visual data, focusing on identifying “what" and “where" are present in a scene, by learning static knowledge mapping from pixels to latent features. In contrast, physics understanding pertains to the ability to simulate dynamic interactions and transitions governed by physical laws, such as motion, collisions, forces, and material behaviors, addressing the “how" and “why" of visual data. In this sense, physics understanding is characterized by its dynamic and predictive nature, which explains why temporal modeling of videos is especially important in this domain. Physics-aware generation aims to combine both understandings to produce content that is both semantically meaningful and physically accurate.
Geometry Understanding vs. Physical Understanding. Geometry understanding involves perceiving and modeling the extrinsic structural properties of objects and scenes. This includes aspects such as shape, size, pose, depth, and positions[208, 209, 210, 211, 212]. Physics understanding focuses on modeling the intrinsic physical properties and behaviors of objects and environments. This involves reasoning about how objects move, deform, and interact under physical laws[93]. In this sense, physics understanding is characterized by its causality and interaction modeling. Integrating both geometry and physics understanding is essential for creating embodied models that not only perceive but also understand and interact with the physical world.
@@ 8 Future Directions
Better Evaluation of Physics-Awareness. Although standardized benchmarks tailored to physical awareness have been introduced recently[183], the evaluation remains ad-hoc or largely relies on manual assessment. The evaluation of physical consistency and dynamic realism in generative models remains a significant challenge that hinders progress. One potential direction is to integrate generative models with established physics engines to evaluate physical plausibility, by converting and deploying generated content within simulation platforms. Incorporating task-oriented evaluations can also be beneficial to assessing physical reasoning capabilities. For example, assessing how well generated content performs in downstream tasks—such as object manipulation in robotics or environmental modeling in autonomous driving—can offer feedback on a model’s utility and physical validity. It is also promising to explore the alignment between generated content and the physical world using embodied agents as an evaluation measurement, for example, by using the behavior similarity of embodied agents running in the simulated and the real world.
Explainability by Physics-Awareness. The integration of physics-aware mechanisms into models holds immense potential for enhancing model explainability. Incorporating explicit physical laws into generative models can offer interpretable pathways linking inputs to outputs, enabling users to trace how specific forces, constraints, and interactions govern the evolution of generated content. By mapping model decisions to well-understood physical principles, the generation process becomes more transparent and aligned with human intuition. Combining physics-aware generative models with existing explainable AI methods, such as saliency maps[213] and feature attribution[214], can further enhance interpretability. For example, visualizing how physical parameters influence specific regions of generated outputs can reveal the model’s reliance on physical properties during generation.
@@
Neural-Symbolic Hybrid Models. The integration of neural and symbolic reasoning presents a promising avenue for advancing physics-aware generation, where symbolic systems can offer rigorous reasoning based on structured knowledge of physics and causality[228, 229]. Embedding symbolic physical constraints, such as differentiable temporal logic[230], can guide generative models to produce outputs that are physically valid and temporally consistent. Symbolic representations also excel in compositional reasoning, allowing systems to understand complex entities and relationships. By incorporating symbolic graphs or ontologies, generative models can achieve more complex compositional generation.
Generative Simulation Engine. Generative simulation engine offers the potential to generate physical simulation directly from textual or semantic prompts, with Genesis[125] as a recent example. Such text-to-simulation models aim to convert high-level textual descriptions into fully interactive, physics-consistent virtual environments. For instance, a command such as “simulate a landslide on a steep mountain after heavy rain" could trigger the generation of a realistic 3D terrain, modeled with appropriate soil and fluid dynamics, and simulate the resulting landslide using physics-based solvers. Generative simulation engines can automate and speed up the design of physical simulation, which can be further integrated into larger systems for physics-aware generation in various modalities.
@@
Applications in Interdisciplinary Domains. Physics-aware generative AI also holds transformative potential across many application domains. Physics-aware generative models can improve climate modeling by enhancing the fidelity, resolution, and scale of simulations[234]. In healthcare, physics-aware generation can significantly enhance surgical training and planning, such as building surgical simulations that accurately replicate the physical properties of tissues to provide realistic training environments[235, 236].
@@ 9 Conclusion
This survey comprehensively reviewed the rapidly evolving field of physics-aware generative AI in vision, highlighting the efforts made to enhance the physical realism and functionality of generated visual content. The core of this survey focused on the integration of physical simulation into generative models, delineating between physics-aware generation with explicit simulation and without explicit simulation. Physics-aware generation stands at a transformative juncture, bridging the gap between virtual and physical realities. We hope that this survey serves as a helpful resource and inspiration for future research in the field of physics-aware generation.

### v2

@@ 6 Physics Evaluation
In this section, we introduce the benchmarks and metrics designed for evaluating the physics awareness of image or video generation models. Conventionally, most of the image and video generation methods are evaluated based on Frechet Inception Distance (FID) [193], CLIP Similarity [194], CLIP-FID [195], and Inception Score (IS) [193]. However, there are a number of gaps in these metrics since they are biased toward the alignment in terms of visual contents and textual semantics. For video generation, the detection of unrealistic patterns violating physical laws is more important, which cannot be detected by these metrics effectively. Furthermore, the evaluation of Fréchet Video Distance (FVD) requires a reference video dataset, which is difficult to obtain for newly generated scenes. These challenges prevent these metrics from effectively evaluating physical commonsense. Some works use Vision Language Models (VLMs) as general evaluators [196, 197], yet these works focus on the evaluation of spatial relationships. Most of them face challenges when generalizing to physical correctness.
@@ 6.1 Benchmarks
6.1.1 Text-Conditioned Benchmarks
Text-conditioned benchmarks evaluate the generated images or videos based on the physical phenomenon described in the prompts.
@@ 6.1.1 Text-Conditioned Benchmarks
PhyBench. PhyBench [198] is a benchmark specifically designed to evaluate the physical commonsense reasoning capabilities of text-to-image models. It defines physical commonsense as four types, including mechanics, optics, thermal,, and material properties. It comprises 700 prompts spanning 31 physical scenarios, where the benchmark is constructed by extending designed prompts describing different types of physical scenarios. For example, given the “pressure" topic, the authors design a prompt “an empty plastic bottle at the bottom of the sea". After that, “empty plastic bottle" will be replaced by other objects, and the scenario “at the bottom of the sea" is also replaced by other scenarios to describe other topics. PhyBench introduces an automated evaluation framework called PhyEvaler, which uses GPT-4o to score generated images based on both scene accuracy and physical correctness.
PhyGenBench. PhyGenBench [199] is a benchmark designed to evaluate the physical commonsense capabilities of text-to-video generation models. It includes 160 carefully crafted prompts that reflect 27 physical laws across the same domains as in PhyBench [198], i.e., mechanics, optics, thermal, and material properties. To support automatic assessment, the authors propose PhyGenEval, a hierarchical evaluation framework that combines vision-language models and GPT-4o for scoring in terms of semantic alignment and physical commensense alignment. PhyGenBench constructs the benchmark for video generation through five steps: (1) Conceptualization: Identify key physical commonsense from four types of physical commonsense. (2) Prompt Engineering: Craft initial prompts to describe the underlying physical phenomenon. (3) Prompt Augmentation: Add additional details to the prompts. (4) Diversity Enhancement: Use GPT-4o to perform object substitution on the augmented prompts. (5) Quality Control: Review the prompts and associated physical laws for accuracy and relevance.
VideoPhy. VideoPhy[200] categorizes the physical interaction types into three categories which are Solid-Solid, Solid-Fluid and Fluid-Fluid. An example of Solid-Solid is the prompt “Bottle topples off the table," where the interaction between “table" and “bottle" is solid-solid, or “Water flows down a circular drain" is the interaction between “water" and “drain" describing Solid-Fluid interaction. Similarly, the prompt “Rain splashing on a pond" is a Fluid-Fluid interaction. Based on these definitions, VideoPhy constructs a total of 688 captions for the benchmark. Through extensive human evaluation across 12 models, including both open (e.g., CogVideoX, OpenSora) and closed (e.g., Pika, Gen-2) models, the authors find that most text-to-video models fail to generate physically coherent outputs. To enable scalable evaluation, the authors propose VideoCon-Physics, a fine-tuned video-language model judging physical plausibility and semantic alignment.
VideoPhy2. VideoPhy2 [201] advances the scope of physical commonsense evaluation in text-to-video generation compared to VideoPhy. VideoPhy2 shifts toward action-centric evaluation, covering 197 diverse real-world actions, including sports, physical activities, and object manipulations, across 3940 prompts. The new dataset introduces longer, more descriptive prompts and augments them with fine-grained annotations of specific physical rules and laws, such as gravity, conservation of mass, and momentum. VideoPhy2 employs a 5-point Likert scale for human evaluations. It also offers an enhanced automatic evaluator, VideoPhy2-AutoEval, which demonstrates alignment with human judgments.
WISA-32K. WISA-32K [20] is a curated video dataset designed to enhance the physical understanding of text-to-video generation models. It contains 32,000 video clips that depict 17 distinct physical phenomena across three core physics domains, i.e., dynamics, thermodynamics, and optics. Each video is annotated with structured physical information, including a textual physical description explaining the underlying principles, qualitative physics categories identifying the types of phenomena present, and quantitative properties such as density, temperature, and time ranges. These annotations are automatically derived using GPT-4o mini and Qwen2-VL. The dataset is constructed through video collection, scene detection, aesthetic filtering, and structured annotation. Experimental results demonstrate that models trained with WISA-32K outperform those trained on general datasets in producing videos that align more closely with real-world physical laws.
6.1.2 Visual-Conditioned Benchmarks
Different from text-conditioned benchmarks, visual-conditioned benchmarks involve the evaluation of the predictions of next frames based on visual comparision of physical properties.
Physics-IQ. Physics-IQ [202] is a visual-conditioned benchmark explicitly designed to evaluate the extent to which video generative models understand physical principles through video prediction. Unlike benchmarks that rely on textual descriptions or synthetic data, Physics-IQ leverages a real-world, high-resolution dataset composed of 396 videos filmed under controlled conditions. Each video depicts physical interactions governed by principles from solid mechanics, fluid dynamics, thermodynamics, optics, and magnetism. The evaluation task is visual-based, where the models are given an initial conditioning segment and predict the next 5 seconds of the sequence. The outputs are then directly compared to ground-truth future frames using spatial, temporal, and pixel-based metrics. These metrics assess not only where and when motion occurs, but also how much motion is present and how accurately it matches the physical trajectory in the real scene.
PisaBench. PisaBench [179] is a benchmark introduced to evaluate the physical accuracy of video generative models through the specific task of modeling object freefall. The benchmark challenges models to generate videos that realistically depict an object falling and potentially colliding with other objects, given only a single input image of the object in midair. The authors curate a dataset composed of 361 real-world slow-motion videos and a set of simulated videos generated using the Kubric engine. PisaBench includes three core evaluations, i.e., using trajectory L2, Chamfer Distance, and Intersection over Union (IoU), which respectively measure the accuracy of object motion, shape fidelity, and spatial consistency over time.
PhyCoBench. PhyCoBench [203] is a benchmark evaluating the physical coherence of text-to-video generation models. The benchmark includes 120 carefully crafted prompts that reflect seven fundamental categories of physical phenomena, including gravity, collision, vibration, friction, fluid dynamics, projectile motion, and rotation. These prompts are inspired by physics textbooks and action recognition datasets, with comprehensive coverage of observable motion scenarios. The authors evaluated four state-of-the-art models through manual rankings based on how realistically the generated videos captured physical behavior. PhyCoBench also introduces PhyCoPredictor, an automated optical flow-guided frame prediction model that uses the first video frame and the associated prompt to forecast motions and video frames. By comparing predictions with the outputs from text-to-video models, PhyCoPredictor assesses physical plausibility using a score based on optical flow and frame reconstruction errors.
6.1.3 Physical Commonsense Ontology
One of the main challenges in the evaluation is the inconsistency in defining physical commonsense. Different lines of works have different ways of definition resulting in difficulties for benchmarking between models. This is one of the key questions that the community should focus on in this field.
COSMOS-Reason1 [191] provides an ontology dividing the physical commonsense into three categories, which are space, time and fundamental physics, with sub-categories in Table III. To evaluate the understanding about physical commonsense, COSMOS-Reason1 [191] collected 5737 questions which have 2828 binary questions and 2909 multiple-choice questions. Among these questions, a subset of 604 questions are associated with 426 video are manually selected to represent physical commonsense benchmark in which 336 are binary questions and 268 are multiple choice questions.
Category Sub-Category
Space Relationship, Plausibility, Affordance, Environment
Time Actions, Order, Causality, Camera, Planning
Fundamental Physics Attributes, States, Object Permanence, Mechanics,
Electromagnetism, Thermodynamics, Anti-Physics
TABLE III: Taxonomy of physical commonsense in COSMOS-Reason1 [191].
@@ 6.2 Evaluation metrics
There are three main approaches for the evaluation metrics of physical awareness. The first approach is based on human evaluation, where humans will give the score to the image or video in terms of physical commonsense with defined criteria. The second way is based on the perception of VLMs. The third way is to use an automatic quantitative score.
Human Evaluation. Human-based evaluation is the most direct and reliable method for assessing the physical plausibility of generated visual content. In this setup, human annotators are asked to judge whether a generated image or video adheres to the laws of physics, matches a given prompt, or displays coherent dynamics. Evaluations can be conducted via rating scales (e.g., binary, three points, or five points) or rankings across different models. For instance, in PhyCoBench [203], humans rank outputs from text-to-video models for each prompt according to the physical realism. PhyGenBench [199] uses manual evaluation that includes semantic alignment and physical commonsense alignment, while VideoPhy [200] and VideoPhy2 [201] introduce human annotations on semantic adherence, physical commonsense, and the grounding of physical rules. Though costly and time-consuming, human evaluation remains essential for establishing reliable ground truth.
VLM-Based Evaluation. Evaluations based on vision-language models automate human judgment by leveraging large multimodal models to assess physical plausibility through language-guided reasoning. These models are either prompted or finetuned to analyze visual content in context of textual descriptions or physics-related questions. For example, PhyBench [198] uses GPT-4o in its PhyEvaler framework to provide item-specific judgments about scene correctness and physical commonsense, mimicking human assessment. PhyGenBench [199] takes this further with PhyGenEval, a multi-stage framework where VLMs detect key physical events, verify temporal order, and assess realism through guided questions and frame comparisons. VideoPhy [200] and VideoPhy2 [201] finetune vision language models using human annotations to predict semantic adherence score, physical commonsense score, and physical rule classification. These VLM-driven approaches offer scalable and cost-efficient alternatives to manual scoring, though their performance still depends on the quality and specificity of prompts or fine-tuning. Moreover, the physical understanding ability of VLMs themselves is still an open question.
Automatic Quantitative Evaluation. This category involves computational metrics that analyze visual content, focusing on low-level or structured measurements of motion, shape, or physical consistency. These metrics are particularly useful for measuring the coherence and predictability of dynamic processes in video prediction. For instance, PhyCoBench [203] introduces PhyCoPredictor, which uses optical flow predictions to detect physical anomalies by comparing predicted and generated motion fields. PISA [179] employs trajectory tracking (L2 distance), shape matching (Chamfer Distance), and object permanence (IoU) to evaluate whether falling object simulations align with expectations. Physics-IQ [202] defines a suite of physical understanding metrics, including Spatial and Spatiotemporal IoU, Weighted Spatial IoU, and Mean Squared Error (MSE), which quantify where, when, and how much motion occurs in generated videos compared to real-world references. These metrics are precise, reproducible, and ideal for benchmarking, though they usually require paried real-world videos as the ground truth.
@@ 7 Discussions
Semantic Awareness vs. Physical Awareness. Semantic awareness and physical awareness serve distinct but complementary roles in how models interpret and generate visual data. Semantic awareness involves interpreting and recognizing concepts and objects within visual data, focusing on identifying “what" and “where" are present in a scene, by learning static knowledge mapping from pixels to latent features. In contrast, physical awareness pertains to the ability to simulate dynamic interactions and transitions governed by physical laws, such as motion, collisions, forces, and material behaviors, addressing the “how" and “why" of visual data. In this sense, physics understanding is characterized by its dynamic and predictive nature, which explains why temporal modeling of videos is especially important in this domain. Physics-aware generation aims to combine both aspects to produce content that is both semantically meaningful and physically accurate.
Geometry Awareness vs. Physical Awareness. Geometry awareness involves perceiving and modeling the extrinsic structural properties of objects and scenes. This includes aspects such as shape, size, pose, depth, and positions[204, 205, 206, 207, 208]. Physical awareness focuses on modeling the intrinsic properties and behaviors of objects and environments. This involves reasoning about how objects move, deform, and interact under physical laws[131]. In this sense, physical awareness is characterized by its causality and interaction modeling. Integrating both geometry and physics is essential for creating embodied models that not only perceive but also understand and interact with the physical world.
@@ 8 Future Directions
Better Evaluation of Physical Awareness. Although standardized benchmarks tailored to physical awareness have been introduced recently[189], the evaluation remains ad-hoc or largely relies on manual assessment. The evaluation of physical consistency and dynamic realism in generative models remains a significant challenge that hinders progress. One potential direction is to integrate generative models with established physics engines to evaluate physical plausibility, by converting and deploying generated content within simulation platforms. Incorporating task-oriented evaluations can also be beneficial to assessing physical reasoning capabilities. For example, assessing how well generated content performs in downstream tasks, such as object manipulation in robotics or environmental modeling in autonomous driving, can offer feedback on a model’s utility and physical validity. It is also promising to explore the alignment between generated content and the physical world using embodied agents as an evaluation measurement, for example, by using the behavior similarity of embodied agents running in the simulated and the real world.
Explainability by Physical Awareness. The integration of physics-aware mechanisms into models holds immense potential for enhancing model explainability. Incorporating explicit physical laws into generative models can offer interpretable pathways linking inputs to outputs, enabling users to trace how specific forces, constraints, and interactions govern the evolution of generated content. By mapping model decisions to well-understood physical principles, the generation process becomes more transparent and aligned with human intuition. Combining physics-aware generative models with existing interpretability methods, such as saliency maps[209] and feature attribution[210], can further enhance explainability. For example, visualizing how physical parameters influence specific regions of generated outputs can reveal the model’s reliance on physical properties during generation.
@@
Neural-Symbolic Hybrid Models. The integration of neural and symbolic reasoning presents a promising direction for advancing physics-aware generation, where symbolic systems can offer rigorous reasoning based on structured knowledge of physics and causality[224, 225]. Embedding symbolic physical constraints, such as differentiable temporal logic[226], can guide generative models to produce outputs that are physically valid and temporally consistent. Symbolic representations also excel in compositional reasoning, allowing systems to understand complex entities and relationships. By incorporating symbolic graphs or ontologies, generative models can achieve more complex compositional generation.
Generative Simulation Engine. Generative simulation engine offers the potential to generate physical simulation directly from textual or semantic prompts, with Genesis[227] as a potential example. Such text-to-simulation models aim to convert high-level textual descriptions into fully interactive, physics-consistent virtual environments. For instance, a command such as “simulate a landslide on a steep mountain after heavy rain" could trigger the generation of a realistic 3D terrain, modeled with appropriate soil and fluid dynamics, and simulate the resulting landslide using physics-based solvers. Generative simulation engines can automate and speed up the design of physical simulation, which can be further integrated into larger systems for physics-aware generation in various modalities. The current main obstacle for generative simulation is its limited generalizability across diverse text prompts and scenes.
@@
Applications in Interdisciplinary Domains. Physics-aware generation also holds transformative potential across many application domains. It can improve climate modeling by enhancing the fidelity, resolution, and scale of simulations[231]. In healthcare, physics-aware generation can significantly enhance surgical training and planning, such as building surgical simulations that accurately replicate the physical properties of tissues to provide realistic training environments[232, 233].
@@ 9 Conclusion
This survey comprehensively reviewed the rapidly evolving field of physics-aware generation in computer vision, highlighting the efforts made to enhance the physical realism and functionality of generated visual content. The core of this survey focused on the integration of physical simulation into generative models, delineating between physics-aware generation with explicit simulation and without explicit simulation. Physics-aware generation currently stands at a transformative juncture, bridging the gap between virtual and physical realities. We hope that this survey serves as a helpful resource and inspiration for future research in the field.